OpenStudy (anonymous):

what is the first quartile of the data set below? 275,257,301,218,265,242,201

OpenStudy (ajprincess):

Arrange the data in ascending order. then (n+1)/4th data is the first quartile. here n is the number of data given

OpenStudy (ajprincess):

what is n+1/4? here n=7

OpenStudy (ajprincess):

guess u r missing 242 in ur arranged list

OpenStudy (anonymous):

yea i just saw that...u wana see my answer choices because theyy are not matching up with what u are saying

OpenStudy (ajprincess):

sure

OpenStudy (anonymous):

201 230 257 218

OpenStudy (ajprincess):

after arranging the data in ascending order u will get it like this 201,218,242,257,265,275,301. let us find n+1/4. n-number of data given. in this case n=7 so n+1/4=7+1/4=2 so the first quartile is the second value in the arranged list. Does tht help? @pookiieluvzyhu

OpenStudy (anonymous):

okay so 218

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

omg thanks so much i got it

OpenStudy (ajprincess):

welcome:)